# DyMO: Training-Free Diffusion Model Alignment with Dynamic Multi-Objective Scheduling

### v2

@@ Abstract
11footnotetext: Corresponding author.
@@ 1 Introduction
Training-free methods align generated images with specific objectives by applying differentiable rewards to adjust the generation process of a pre-trained diffusion model during inference [53, 51, 40, 54]. To enhance text-image alignment, classifier guidance [8, 38] uses the gradient of the pre-trained image classifier. The idea is extended to alignment with different types of conditional signals [53, 51, 45], and some training-free methods [43, 40, 7] also consider aesthetic improvement through optimizing related objectives. Unlike training-based methods, most training-free methods are mainly restricted to the alignment of specific conditions (e.g., class). To achieve more general applications, in this paper, we consider the training-free alignment of general preference in inference. The learned text-aware human preference scores [19, 46, 23] measure both text-image alignment and human visual preferences. We can perform training-free alignment with the differentiable scores in score-based diffusion models (SBDMs) [38, 37]. Considering that the preference scores perform on clean images by default, specific designs are required to incorporate the guidance to the denoising process in diffusion models. It needs retraining of the noise-aware or step-dependent score estimator [17, 27, 25, 2] or full-chain backpropagation from the output to intermediate steps [40, 43, 7], which is data or time demanding or time expensive, respectively. For memory and time efficiency, we use one-step clean image approximation from noisy image [53, 35, 51], which causes underestimated guidance. Especially, compared to the visual characteristics, the semantic contents (e.g., entities and layout) tend to be determined in early noisy steps. Although the text-aware preference scores can reflect the text-image semantic alignment, it is more challenging the guide the semantic contents due to the blurred predicted image from noisy samples.
@@ 2.1 Diffusion Model Fine-Tuning
Recently, diffusion models have rapidly advanced in image generation, with various methods emerging to improve quality and align with human preference. One simple yet effective way [29, 48, 6, 47, 9] is to use differentiable reward functions as the objective, enabling model optimization through gradient backpropagation. To tackle the limitations of non-differentiable reward functions and unstable training dynamics, some RL-based methods [21, 4, 12] are proposed for for better image-text alignment. Following the success of Direct Preference Optimization (DPO) [30] by eliminating the need for explicit reward models, D3PO [49] fine-tuned the model on the preferred and dispreferred image pairs based on human evaluators. Diffusion-DPO [44] re-formulated DPO for model optimization on the preference dataset[19]. SPO [23] further generalize d DPO paradigm at each denoising step to ensure accurate step-aware preference alignment. DenseReward [50] introduced a temporal discounting factor to prioritize early steps for better preference alignment. NCPPO [13] emphasized preference optimization in perceptual feature space with more information. Diffusion-RPO [14] made contrastive weights for image pairs and Diffusion-KTO [22] defined the alignment objective as human utility maximization. Despite notable progress in diffusion model alignment, discrepancies persist between model performance and human preferences, along with large computational demands.
@@ 4 Methodology
We propose a training-free framework to align with contextual semantics and human preferences. Specifically, we utilize the preference loss $\mathcal{L}_{R}(\mathbf{x}_{0|t}^{\prime},\mathbf{c},t)$ to estimate the text-aware human preference score of denoised image $\mathbf{x}_{0|t}^{\prime}$ generated by decoding the one-step predicted latent feature $\mathbf{z}_{0|t}^{\prime}$ from $\mathbf{z}_{t}$. Then we compute the gradient of the reward score to guide the denoising process like Eq. 6. However, the predicted images from noisy samples in early steps are highly blurred, resulting in the inaccurate estimation of $\mathcal{L}_{R}$. We dive into the cross attention mechanism and discover rich content semantics in early steps. We propose a semantic align objective (Sec. 4.1), using to minimize the discrepancies between image content semantics and text semantics extracted via LLMs. We design the tailored guidance via dynamically scheduling two objectives (Sec. 4.2), generating detailed content while preserving the layout. Besides, we propose a dynamic time-travel strategy to offer better guidance for diffusion models (Sec. 4.3). The whole pipeline is demonstrated in Fig. 2 and Algorithm 1.
@@ 4.1 Semantic Alignment Guidance
Attention map exploration. During the initial chaotic stage, the basic image structure takes shape quietly. However, the sample $z_{t}$ exhibits a high level of noise, causing that the approximately predicted image $x_{0|t}^{\prime}$ is extremely blurred. Consequently, training-free guidance is hard to make anything reasonable, providing ineffective semantic supervision. Previous works [54, 16, 42] show attention maps reflect the semantic contents, $M=QK^{T}/{\sqrt{d}}$, where $d$ is the feature dimension, queries $Q$ and keys $K$ are the linear projection of the intermediate image feature and text embedding of prompt $\mathbf{c}$, respectively. We denote $M_{u}$ as the cross attention map for the $u$-th word token within the prompt. As shown in Fig. 2, cross attention maps contain rich content semantics, which can be used to align with semantics of text prompt for early supervision.
Semantic capture from text prompts. Upon input of a complex scene prompt $\mathbf{c}$ from the user, we leverage LLMs (e.g., GPT-4 [1], Llama [41]) to envision the words likely to be generated in the final image and capture the entities and their corresponding attributes: $\{E_{i}\}_{i=1}^{n},\{A_{i}\}_{i=1}^{n}$, where $n$ is the number of the entities identified in the given prompt using LLM and $E_{i}$ represents $i$-th entity. Besides, $A_{i}$ denotes the attribute set associated with $E_{i}$, encompassing features like color, shape, texture, etc. It is noted that each entity possesses a unique set of attributes in varying quantities, which are determined dynamically through LLMs-driven heuristics.
Text semantic graph construction. In order to achieve semantic alignment, we construct a semantic graph $\mathbf{G}=(\mathbf{N},\mathbf{S})$ to find the internal relationships within the prompts for manipulating attention maps, where $\mathbf{N}=\{E_{i}\}_{i=0}^{n}\cup\{A_{i}\}_{i=0}^{n}$ is the set of nodes in the graph $\mathbf{G}$. For the set of edges $\mathbf{S}$, we propose two kinds of criteria to determine all possible relationships. We hypothesize that each entity exists independently within the image, occupying its own distinct position and space. Consequently, the relationships between any two entities are negative. In addition, correct attribute binding is essential to enhance image-prompt alignment. The relationships between the entities and their corresponding attributes are positive:
$\mathbf{S}_{\text{pos}}=\bigg{\{}\sum_{i=1}^{n}\sum_{j=1}^{\|A_{i}\|}(E_{i},A_{ij})\bigg{\}},$ (7)
@@
where $\mathbf{S}_{\text{pos}}$ and $\mathbf{S}_{\text{neg}}$ represent positive and negative pair sets, respectively. $A_{ij}$ denotes the $j$-th attirbute of the $i$-th entity. The set of edges is obtained by $\mathbf{S}=\mathbf{S}_{\text{pos}}\cup\mathbf{S}_{\text{neg}}$.
Semantic alignment objective. Based on the text semantic graph, we design a semantic alignment loss to achieve better layout composition and precise attribute binding in early steps:
$\mathcal{L}_{A}=$ $-\frac{1}{\left|\mathbf{S}_{\text{pos}}\right|}\sum_{(s,l)\in\mathbf{S}_{\text{pos}}}f(M_{s},M_{l})$ (9)
$+\frac{1}{\left|\mathbf{S}_{\text{neg}}\right|}\sum_{(s,l)\in\mathbf{S}_{\text{neg}}}f(M_{s},M_{l}),$
where $(s,l)$ denotes the edges in $\mathbf{S}$ and $f(\cdot)$ is the cosine similarity function. The proposed layout loss mitigates the catastrophic object neglect problem by isolating each entity, which prevents the overlap among different objects and helps reduce semantic information loss. Meanwhile, it enhances accurate attribute association by encouraging higher alignment of cross attention maps between objects and related attributes.
@@
1: Input: prompt $\mathbf{c}$, noise predictor $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\cdot,t)$, human preference evaluator $\mathcal{L}_{R}(\cdot,\mathbf{c})$, layout loss function $\mathcal{L}_{A}(\cdot)$, timesteps $T$, decoder $D$, guidance strength $\eta_{t}$ and pre-defined parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $h_{t}$, $k$.
@@ 4.2 Multi-Objective Dynamic Scheduling
Preference alignment. After the overall layout of the generated image is roughly established, we incorporate human feedback to guide conditional image generation, aligning the output with human preference in intermediate generation steps. Specifically, we adopt a pre-trained preference model as the reward function $\mathcal{L}_{R}$ to update the $\mathbf{z}_{t-1}$ as Eq. 6. However, directly applying the preference loss can unintentionally alter the layout because the predicted image $\mathbf{x}_{0|t}^{\prime}$ remains blurred, which leads to inaccurate feedback for the preference model. To address this issue, we propose a solution by adaptively combining the layout loss and preference loss:
@@ 4.3 Improving Training-free Guidance
Dynamic Time-Travel Straregy. After rethinking the efficient time-travel strategy from FreeDom [53], we discover the fixed iteration steps hinder convergence efficiency since the abilities of each denosing step in diffusion models are different. We propose dynamic recurrent scheme to dynamically schedule the number of iterations at each denosing step in diffusion models are different. We compute the iteration count $r_{t}=h_{t}\cdot\left\|\boldsymbol{g}_{t}\right\|$ only once at the start of each step where $h_{t}$ is a pre-defined parameter. When $\left\|\boldsymbol{g}_{t}\right\|$ is larger, more iterations are required to promote convergence, and vice versa.
@@
5 Experiments
In this section, different experiments are conducted on a wide range of generative models to verify the effectiveness and flexibility of our method. We first describe the foundational setup for the experiments and the configuration of the parameters setting. Then we evaluate the performance of our method with various generative backbones, and make comprehensive comparison with state-of-the-art approaches. Besides, we delve deeper into the role of each component for the guidance during the denoising process, supported by ablation studies for further analysis.
5.1 Experimental Setting
Implementation Details. Our method requires no training and is compatible with various generative models. We employ the GPT-4 [1] as the base LLMs to construct semantic graph. Additionally, we adopt the pre-trained step-aware preference model from SPO [23] to facilitate human feedback, which is developed from PickScore [19] model and fine-tuned on the human preference pairs dataset, Pick-a-Pic [19]. All our experiments are achieved using an NVIDIA A100 GPU for SDXL-based methods or V100 GPU for SD V1.5-based methods.
Datasets and Metrics. We utilized three datasets to validate the effectiveness of our method: 500 unique prompts of the Pick-a-Pic validation set, 500 prompts from HPSv2 benchmark [46] and 1000 prompts from Partiprompt [52]. We choose four AI feedback models to assess the image quality: PickScore [19] (general human preference), HPSv2 [46] (prompt alignment), ImageReward [48] (general human preference), Aesthetic Predictor [34] (non-text-based visual appeal). For all metrics, higher values indicate better performance.
5.2 Comparison with SOTA Methods
To verify the effectiveness of our approach, we compare our method with training-based and training-free models. The former comprises human preference learning methods, including AlignProp [29], Diffusion-DPO [44], Diffusion-KTO [22], SPO [29]. The latter includes DNO [40], PromptOpt [7], FreeDom [53]. For SDXL-based frameworks, we further make a comparison with the latest models with advanced network architectures, such as FLUX [20], SD V3.5 [31].
@@ 5.2 Comparison with SOTA Methods
Figure 3: Qualitative comparison based on SD V1.5 backbones. The text prompts used to generate the images are provided in the Tab. 6.
@@ 5.2.1 Quantitative Analysis
To objectively evaluate the performance of our method, we conduct a quantitative comparison on the Pick-a-Pic dataset, with the results organized in Tab. 1 for SD V1.5-based backbones and Tab. 2 for SDXL-based backbones, respectively. Under the guidance of our method, models including SD V1.5 [31], SDXL [28], Diffusion-DPO [44], and SPO [23] achieve notable improvements, surpassing even advanced models like FLUX [20] and SD V3.5 [31]. Specifically, the training-free methods (DNO [40], PromptOpt [7], FreeDom [53]) suffer from underestimation in early steps, resulting in the inferior outcomes. AlignProp [29] introduces the dependency bias via truncated backpropagation. Diffusion-DPO [44] and Diffusion-KTO [22] produce suboptimal results due to trajectory-level preference supervision, causing inaccurate alignment. SPO [23] lose the layout and semantic information in early denoising steps since step-wise preference model struggles to provide reliable preference. In contrast, our training-free method provides tailored guidance for different denoising steps, generating high-quality images with better layout composition and human preference alignment.

### v3

@@ Abstract
11footnotetext: D. Gong is the corresponding author. This project was partially supported by an ARC DECRA Fellowship (DE230101591) to D. Gong.
@@ 1 Introduction
Training-free methods align generated images with specific objectives by applying differentiable rewards to adjust the generation process of a pre-trained diffusion model during inference [54, 52, 41, 55]. To enhance text-image alignment, classifier guidance [8, 39] uses the gradient of the pre-trained image classifier. The idea is extended to alignment with different types of conditional signals [54, 52, 46], and some training-free methods [44, 41, 7] also consider aesthetic improvement through optimizing related objectives. Unlike training-based methods, most training-free methods are mainly restricted to the alignment of specific conditions (e.g., class). To achieve more general applications, in this paper, we consider the training-free alignment of general preference in inference. The learned text-aware human preference scores [20, 47, 24] measure both text-image alignment and human visual preferences. We can perform training-free alignment with the differentiable scores in score-based diffusion models (SBDMs) [39, 38]. Considering that the preference scores perform on clean images by default, specific designs are required to incorporate the guidance to the denoising process in diffusion models. It needs retraining noise-aware or step-dependent score estimator [18, 28, 26, 2] or full-chain backpropagation from the output to intermediate steps [41, 44, 7], which is data or time demanding or time expensive, respectively. For memory and time efficiency, we use one-step clean image approximation from noisy image [54, 36, 52], causing underestimated guidance. Especially, compared to the visual characteristics, the semantic contents (e.g., entities and layout) tend to be determined in early noisy steps. Although the text-aware preference scores can reflect the text-image semantic alignment, it is more challenging to guide the contents due to the blurred predicted image from noisy samples.
@@ 2.1 Diffusion Model Fine-Tuning
Recently, diffusion models have rapidly advanced in image generation, with various methods emerging to improve quality and align with human preference. One simple yet effective way [30, 49, 6, 48, 9] is to use differentiable reward functions as the objective, enabling model optimization through gradient backpropagation. To tackle the limitations of non-differentiable reward functions and unstable training dynamics, some RL-based methods [22, 4, 12] are proposed for better image-text alignment. Following the success of Direct Preference Optimization (DPO) [31] by eliminating the need for explicit reward models, D3PO [50] fine-tuned the model on the preferred and dispreferred image pairs based on human evaluators. Diffusion-DPO [45] re-formulated DPO for model optimization on the preference dataset[20]. SPO [24] further generalized DPO paradigm at each denoising step to ensure accurate step-aware preference alignment. DenseReward [51] introduced a temporal discounting factor to prioritize early steps for better preference alignment. NCPPO [13] emphasized preference optimization in perceptual feature space with more information. Diffusion-RPO [15] made contrastive weights for image pairs and Diffusion-KTO [23] defined the alignment objective as human utility maximization. Despite notable progress in diffusion model alignment, discrepancies persist between model performance and human preferences, along with large computational demands.
@@ 4 Methodology
Overview. We propose a training-free framework to align generated contents with contextual semantics and human preferences. We utilize a preference loss $\mathcal{L}_{R}(\mathbf{x}_{0|t}^{\prime},\mathbf{c},t)$ to estimate the text-aware human preference score [20, 24, 49] of denoised image $\mathbf{x}_{0|t}^{\prime}$ generated by decoding the one-step predicted latent feature $\mathbf{z}_{0|t}^{\prime}$ from $\mathbf{z}_{t}$. We compute the gradient of the reward score to guide the denoising process like Eq. 6. Estimating $\mathcal{L}_{R}$ and propagating its guidance to early steps is ineffective, where the images are highly noisy or the predicted outputs are highly blurred. To guide image content in the early steps, we leverage text-image attention maps to capture semantic information in the noisy images, as the content is shaped by the attention. We propose a semantic align objective (Sec. 4.1), through minimizing the discrepancies between image content semantics (as reflected in attention maps) and text semantics extracted via LLMs. We design the tailored guidance via dynamically scheduling two objectives (Sec. 4.2), generating detailed content while preserving the layout. Besides, we propose a dynamic time-travel strategy to offer better guidance for diffusion models (Sec. 4.3). The whole pipeline is demonstrated in Fig. 2 and Algorithm 1.
@@ 4.1 Semantic Alignment Guidance
Attention map exploration. During the initial chaotic stage, the basic image structure takes shape quietly. However, the sample $z_{t}$ exhibits a high level of noise, causing that the approximately predicted image $x_{0|t}^{\prime}$ is extremely blurred. Consequently, training-free guidance is hard to make anything reasonable, providing ineffective semantic supervision. Previous works [55, 17, 43] show attention maps reflect the semantic contents, $M=QK^{T}/{\sqrt{d}}$, where $d$ is the feature dimension, queries $Q$ and keys $K$ are the linear projection of the intermediate image feature and text embedding of prompt $\mathbf{c}$, respectively. In T2I Diffusion models, each text token $u$ is injected into the image via cross-attention with an attention map $M_{u}$. We use $M_{u}$ to supervise early-stage generation by aligning image semantics with the text prompt.
Semantic capture from text prompts. Upon input of a complex scene prompt $\mathbf{c}$ from the user, we leverage LLMs (e.g., GPT-4 [1] and Llama [42]) to analyze the words likely to present in the final image and explicitly extract the entities and their corresponding attributes: $\{E_{i}\}_{i=1}^{n},\{A_{i}\}_{i=1}^{n}$, where $n$ is the number of the entities identified in the given prompt using LLM and $E_{i}$ represents $i$-th entity, $A_{i}$ denotes the attribute set associated with entity node $E_{i}$, encompassing features like color, shape, texture, etc. Each entity possesses a unique set of attributes in varying quantities, which are determined dynamically by LLM-driven heuristics.
Text semantic graph construction. In order to achieve semantic alignment, we construct a semantic graph $\mathbf{G}=(\mathbf{N},\mathbf{S})$ to find the internal relationships within the prompts for manipulating attention maps, where $\mathbf{N}=\{E_{i}\}_{i=0}^{n}\cup\{\tilde{A}_{i}\}_{i=0}^{n}$ is the set of nodes in the graph $\mathbf{G}$, where $\tilde{A}_{i}$ denotes the attribute elements in $A_{i}$. For the set of edges $\mathbf{S}$, we propose two kinds of criteria to determine all possible relationships. We hypothesize that each entity exists independently within the image, occupying its own distinct position and space. The relationships between any two entities are negative. The relationships between the entities and their corresponding attributes are positive, for correct binding between entities and attributes:
$\mathbf{S}_{\text{pos}}=\{(E_{i},A_{ij})\mid\forall i=1,\cdots,n,\forall j=1,\cdots,|A_{i}|\},$ (7)
@@
where $\mathbf{S}_{\text{pos}}$ and $\mathbf{S}_{\text{neg}}$ represent positive and negative pair sets, respectively. $A_{ij}$ denotes the $j$-th attirbute of the $i$-th entity, with a slight abuse of notations. The set of edges is obtained by $\mathbf{S}=\mathbf{S}_{\text{pos}}\cup\mathbf{S}_{\text{neg}}$.
Semantic alignment objective. Each entity or attribute corresponds to an attention map, denoted by $M_{u}$. We design a semantic alignment loss to align the image contents ($M_{u}$s’ relationships) with the text semantic graph represented by ($\mathbf{S}_{\text{pos}}$ and $\mathbf{S}_{\text{neg}}$), aiming to achieve better layout composition and precise attribute binding in early steps:
$\mathcal{L}_{A}=$ $-\frac{1}{\left|\mathbf{S}_{\text{pos}}\right|}\sum\nolimits_{(s,l)\in\mathbf{S}_{\text{pos}}}f(M_{s},M_{l})$ (9)
$+\frac{1}{\left|\mathbf{S}_{\text{neg}}\right|}\sum\nolimits_{(s,l)\in\mathbf{S}_{\text{neg}}}f(M_{s},M_{l}),$
where $(s,l)$ denotes the edges in $\mathbf{S}$ and $f(\cdot)$ is the cosine similarity function. If more than one tokens are generated from a single object (e.g. $u_{1},u_{2}$), this case would be detected, and the attention maps $M_{u_{1}},M_{u_{2}}$ would share the correspondences to the same set of attributes. The proposed semantic alignment loss mitigates the catastrophic object neglect problem by isolating each entity, which prevents the overlap among different objects and helps reduce semantic information loss. Meanwhile, it enhances accurate attribute association by encouraging higher alignment of cross attention maps between objects and related attributes.
@@
1: Input: prompt $\mathbf{c}$, noise predictor $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\cdot,t)$, human preference evaluator $\mathcal{L}_{R}(\cdot,\mathbf{c})$, semantic alignment loss function $\mathcal{L}_{A}(\cdot)$, timesteps $T$, decoder $D$, guidance strength $\eta_{t}$ and pre-defined parameters $\beta_{t}$, $\bar{\alpha}_{t}$, $h_{t}$, $k$.
@@ 4.2 Multi-Objective Dynamic Scheduling
Preference alignment. After the overall layout of the generated image is roughly established, we incorporate human feedback to guide conditional image generation, aligning the output with human preference in intermediate generation steps. Specifically, we adopt a pre-trained preference model [20, 24] as the reward function $\mathcal{L}_{R}(\mathbf{x}_{0|t}^{\prime},\mathbf{c},t)=\text{exp}(\tau\cdot f_{\text{V}}(\mathbf{x}_{0|t}^{\prime},t)^{T}f_{\text{T}}(\mathbf{c}))$ to update the $\mathbf{z}_{t-1}$ as Eq. 6, where $\tau$ is the temperature and $f_{\text{V/T}}$ represents vision/text encoder of a CLIP-style model trained on preference data. However, directly applying the preference loss can unintentionally alter the layout because the predicted image $\mathbf{x}_{0|t}^{\prime}$ remains blurred, which leads to inaccurate feedback for the preference model. To address this issue, we propose a solution by adaptively combining the semantic alignment loss and preference alignment loss:
@@ 4.3 Improving Training-free Guidance
Dynamic Time-Travel Straregy. After rethinking the efficient time-travel strategy from FreeDom [54], we discover the fixed iteration steps hinder convergence efficiency since the abilities of each denosing step in diffusion models are different. We propose dynamic recurrent scheme to dynamically schedule the number of iterations at each denosing step in diffusion models are different. We compute the iteration count $r_{t}=h_{t}\cdot\left\|\boldsymbol{g}_{t}\right\|$ only once at the start of each step where $h_{t}$ is a pre-defined parameter. When $\left\|\boldsymbol{g}_{t}\right\|$ is larger, more iterations are required to promote convergence.
@@
Figure 3: Qualitative comparison based on SD V1.5 backbones.
5 Experiments
In this section, different experiments are conducted on a wide range of generative models to verify the effectiveness and flexibility of our method. We validate the performance of our method through comparing with various generative backbones and existing state-of-the-art approaches. We also delve deeper into the role of each component for the guidance during the denoising process, supported by ablation studies for further analysis.
5.1 Experimental Setting
Implementation Details. Our method requires no training and is compatible with various generative models. We employ the GPT-4 [1] as the base LLMs to construct semantic graph. Additionally, we adopt the pre-trained step-aware preference model from SPO [24] to facilitate human feedback, developed from PickScore [20] model and fine-tuned on the human preference pairs dataset, Pick-a-Pic [20]. All our experiments uses an NVIDIA A100 GPU for SDXL-based methods or V100 GPU for SD V1.5-based methods.
Datasets and Metrics. We utilized three datasets to validate the effectiveness of our method: 500 unique prompts of the Pick-a-Pic validation set, 500 prompts from HPSv2 benchmark [47] and 1000 prompts from Partiprompt [53]. We choose four AI feedback models to assess the image quality: PickScore [20] (general human preference), HPSv2 [47] (prompt alignment), ImageReward [49] (general human preference), Aesthetic Predictor [35] (non-text-based visual appeal). For all metrics, higher values indicate better performance. All test images are generated using 50 denoising steps during inference.
5.2 Comparison with Existing Methods
To verify the effectiveness of our approach, we compare our method with training-based and training-free models. The former comprises human preference learning methods, including AlignProp [30], Diffusion-DPO [45], Diffusion-KTO [23], SPO [30]. The latter includes DNO [41], PromptOpt [7], FreeDom [54]. For SDXL-based frameworks, we make a comparison with the latest models with advanced network architectures (FLUX [21] and SD V3.5 [32]).
@@ 5.2.1 Quantitative Analysis
To objectively evaluate the performance of our method, we conduct a quantitative comparison on the Pick-a-Pic dataset, with the results organized in Tab. 1 for SD V1.5-based backbones and Tab. 2 for SDXL-based backbones, respectively. Under the guidance of our method, models including SD V1.5 [32], SDXL [29], Diffusion-DPO [45], and SPO [24] achieve notable improvements, surpassing even advanced models like FLUX [21] and SD V3.5 [32]. Specifically, the training-free methods (DNO [41], PromptOpt [7], FreeDom [54]) suffer from underestimation in early steps, resulting in the inferior outcomes. AlignProp [30] introduces the dependency bias via truncated backpropagation. Diffusion-DPO [45] and Diffusion-KTO [23] produce suboptimal results due to trajectory-level preference supervision, causing inaccurate alignment. SPO [24] lose the layout and semantic information in early steps. In contrast, our training-free method provides tailored guidance for different denoising steps, generating high-quality images with better layout composition and human preference alignment.